Shift report 20000809 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 09 2000 - 00:11:14 EDT

  • Next message: http account: "Shift report 20000809 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Aug  9 00:11:14 2000, Shift: 16:00-24:00
    Supervisor: TT, with: JN
    Log book#4 pages 32 - 37
    
    Shift Summary:
    ------------------------------------------------------------
    1. Power supply work until 20.30, then injection and 
       ramping..
       21.12: Beam stored and cogged at gamma=70.  Steering 
       for Phenix.
       22.12: Attempt to steer for Brahms.
    
    2. Run 2144 from 22.14 to 23.27.  Varying scaler rates 
       (see logbook)
       Contains 477 collision events, probably good run!
       Interaction vertex seems to be at right place.  Tracks 
       seen in all 4 TPCs.
       T2b anode voltage raised from 1250 V to 1280 V before 
       starting run 2145 (still running.)  
       Some improvement in cluster quality.
       
    3. No special problems (except MTPC2 and T2a anode wires 
       tripped once due to sudden beam injection during HV 
       shift check test, and T3 anode wires were found tripped 
       at the same time.)  Otherwise stable.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 09 2000 - 00:14:40 EDT